Calculate molar mass 118.936 g/mol




1
Expert's answer
2021-01-27T04:53:04-0500

A compound with molar mass of 118.936g/mol is permanganate ions (MnO4-).

To calculate its molar mass one should sum the product of atomic mass of each element with its subscript.

Mn=1*54.94=54.94

O4=4*15.999=63.996

"\therefore" Molar mass= 63.996+54.94= 118.936g/mol